OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (2, -3) and is perpendicular to y =2/3 x + 3.

OpenStudy (anonymous):

you take the opposite reciprocal of 2/3 which would be -3/2 then you put all of your into into the point slope form which would be y+3=-3/2(x-2) then you solve for y!

OpenStudy (anonymous):

y=3/2x-6

OpenStudy (anonymous):

-3/2 is the new slope